Compute the numerical value of the Gravimetric factor for Magnesium ammonium phosphate hexahydrate (MgNH4PO4.6H2O) precipitate into nitrogen.


1
Expert's answer
2021-03-22T06:20:41-0400


Gravimetric Factor = Mass of MgNH4PO4.6H2O/ molecular Mass of MgNH4PO4.6H2O +N2

Molar Mass of MgNH4PO4.6H2O = 24.3050 + 14.0067 + 30.973762 + 4(15.994) + 6(2.01588 + 15.994) = 245. 4065 g/mol

Mass of MgNH4PO4.6H2O = 245. 4065 g/mol × 1mol = 245.4065 g

Molecular Mass of MgNH4PO4.6H2O + N2

= 24.3050 + 14.0067 + 30.973762 + 4(15.994) + 6(2.01588 + 15.994) + (14.0066)2 = 273.4199 g

Gravimetric Factor = 245.4065/273.4199

= 0.897544